OpenStudy (aonz):

Simutaneosly solve x-5y-z=13 2x+y+z=0 -3x+y+2z=11

OpenStudy (aonz):

I know how to solve 2 simutaneous equations but no idea about 3

OpenStudy (aonz):

i rather the elimination method :P

hartnn (hartnn):

ok, so if i help you get 2 equations in x and y only , will you be able to solve the rest ? 1) Add 1st and 2nd equation, what u get ?

OpenStudy (aonz):

3x-4y=13

OpenStudy (aonz):

i just need the steps in which to complete it i think

hartnn (hartnn):

now multiply 1st equation by 2

OpenStudy (aonz):

6x-8y=26

hartnn (hartnn):

:O how ?

OpenStudy (aonz):

2x-10y-2z=26

hartnn (hartnn):

now add it with 3rd equation.....

OpenStudy (aonz):

-x-9y=37 Btw does it matter if you add the 2nd to the 3rd eqution instead of 1st to 3rd?

hartnn (hartnn):

if you add 2nd and 3rd equation, you could not have eliminated 'z'

hartnn (hartnn):

now can you find 'x' and 'y' ?

OpenStudy (aonz):

y=-4?

OpenStudy (aonz):

x=-1

hartnn (hartnn):

yes.

hartnn (hartnn):

now can u find z ?

OpenStudy (aonz):

mhm... I sub x and y into the 1,2,3 equation?

OpenStudy (aonz):

ok z =6 thank you

hartnn (hartnn):

yes, correct :) ask if any doubts.